# CLAMP: Crowdsourcing a LArge-scale in-the-wild haptic dataset with an open-source device for Multimodal robot Perception

### v1

@@ B.4 Data annotation
The prompt contains text along with more than 12 in-context examples. The prompt is as follows:
@@ 3. Heterogenous Surfaces:
- Under the Heterogenous Surfaces section, specify if the object is made up of different materials on the two opposing sides
Each in-context example is provided in the following format (the horizontal line demarcates the two inputs sent as one user prompt and the assistant prompt containing the example annotation
@@
1. Object: cup 2. Materials: steel 3. Heterogenous Surfaces: No
@@ B.5 Dataset statistics
We demonstrate the diversity in the CLAMP dataset along three axes: distribution of object materials, grasping forces, and grasping speeds. In Figure 6, we show the spread in values of maximum force, maximum grasping speed, along with object material. Additionally, we plot the duration of contact instances (number of samples in a contact instance before padding) as a histogram.
For the material distribution, we only consider objects that do not have heterogenous surfaces. Our dataset contains 5357 objects, out of which 680 ($\sim 13\%$) feature objects with heterogeneous surfaces. We plot the distribution of material labels for contact instances with the remaining homogenous objects. Figure 6 shows that the material labels contain significant class imbalance, with the largest class size (hard plastic) containing approximately 1000x as many samples as the smallest class (dry wall).
@@ C.1 Model training details
Haptic encoder details. We use the InceptionTime architecture [51], choosing the number of filters as 256. Our model uses 7 1-dimensional convolutional kernels, the smallest and largest ones being of length 7 and 250 respectively, that are applied on each latent feature. These components make up one of the 6 Inception blocks in our model. The 6 blocks are followed by a 3-layer MLP that take a latent embedding of dimension 2048 as input, and returns logits of dimension 14 as output.
@@ C.2 Extracting Visual Features from GPT-4o
To obtain a material prediction and log-probabilities, we employ a two-step prompting approach for better material classification. We initially feed in an image to identify the object in question. Then, we further prompt GPT with the prediction object and input image to generate a prediction for object material. In-context examples are provided for only the second step. For this step, we request the top 14 logits. We filter out logits with tokens that do not belong to the list of materials, and assign a probability of 0 to those material classes for which a log-probability was not generated.
@@ Appendix D Real robot experiments
Table 5: Rules for sorting recyclables
For all real-robot experiments, we assume the grasp of each object. For the sorting and metallic object retrieval experiments, each object also has a pre-defined pose from where an image is captured.
@@ D.1 Feature extraction for parallel jaw gripper embodiments
For impedance, we use the ratio of the force difference to linear acceleration feature. A lower threshold is not needed as the robots grasp objects with pre-defined speeds.
D.2 Sorting recyclable from non-recyclable items
@@
D.3 Retrieving objects from a cluttered bag
For this experiment, we place three objects in a cloth bag: a T-shirt, a steel fork, and a bag of popcorn. We use a total of two sequences across the 13 trials: fork $\rightarrow$ T-shirt $\rightarrow$ popcorn and T-shirt $\rightarrow$ fork $\rightarrow$ popcorn. We observe that grasping the bag first or second causes the cloth bag to crumple and thus prevent further grasps.
D.4 Separating ripe from overripe bananas

### v2

@@ B.4 Data annotation
The prompt contains text along with more than 12 in-context examples containing text and images. The prompt is as follows:
@@ 3. Heterogeneous Surfaces:
- Under the heterogeneous Surfaces section, specify if the object is made up of different materials on the two opposing sides
Each in-context example is provided in the following format (the horizontal line demarcates the two inputs sent as one user prompt and the assistant prompt containing the example annotation)
@@
1. Object: cup 2. Materials: steel 3. heterogeneous Surfaces: No
@@ B.5 Dataset statistics
We demonstrate the diversity in the CLAMP dataset along three axes: distribution of object materials, grasping forces, and grasping speeds. In Figure 6, we show the spread in values of maximum force, maximum grasping speed, along with object material. Additionally, we plot the duration of contact instances (before padding) as a histogram.
For the material distribution, we only consider objects that do not have heterogeneous surfaces. Our dataset contains 5357 objects, out of which 680 ($\sim 13\%$) objects are objects with heterogeneous surfaces. We plot the distribution of material labels for contact instances with the remaining homogeneous objects. Figure 6 shows that the material labels contain significant class imbalance, with the largest class size (hard plastic) containing approximately 1000x as many samples as the smallest class (dry wall).
@@ C.1 Model training details
Haptic encoder details. We use the InceptionTime architecture [31], choosing the number of filters as 256. Our model uses 7 1-dimensional convolutional kernels, the smallest and largest ones being of length 7 and 250 respectively, that are applied on each latent feature. These components make up one of the 6 Inception blocks in our model. The 6 blocks are followed by a 3-layer MLP that takes a latent embedding of dimension 2048 as input, and returns logits of dimension 14 as output.
@@
Loss function used for training CLAMP model. For the KL-divergence loss, we observe empirically that treating the vision distribution as the target and visuo-haptic as the student performs better than vice-versa.
@@ C.2 Extracting Visual Features from GPT-4o
To obtain a material prediction and log-probabilities, we employ a two-step prompting approach for better material classification. We initially feed in an image to identify the object in question. Then, we further prompt GPT with the predicted object and input image to generate a prediction for object material. In-context examples are provided for only the second step. For this step, we request the top 14 logits. We filter out logits with tokens that do not belong to the list of materials, and assign a probability of 0 to those material classes for which a log-probability was not generated.
@@ Appendix D Real robot experiments
Table 5: Rules for trash sorting
@@ D.1 Feature extraction for parallel jaw gripper embodiments
For impedance, we use the ratio of the force difference to linear acceleration feature. A lower threshold is not needed as the robots grasp objects with predefined speeds.
D.2 Data collection for robot embodiments
For all real-robot experiments, we assume the grasp of each object. For the sorting and metallic object retrieval experiments, each object also has a pre-defined pose from where an image is captured. For objects that are unstable (e.g. cello tape), we hold them or weakly attach them to the table using adhesive. This is not dissimilar to human data collection as users report holding the object with one hand while grasping using the CLAMP device in the other hand. Finally, for each object, we capture an image using the CLAMP device camera, to create the dataset.
The dataset for the “Franka with CLAMP device” embodiment is 60 objects, while datasets for the parallel jaw embodiments is 30 objects each. Comparing finetuned model performance across embodiments, we hypothesize that “Franka with CLAMP” underperforms “Franka with Parallel Jaw” because it includes data from wide objects that the parallel jaw embodiment cannot grasp. Because the robot grasping policy was open-loop, this led to variable and noisy contact between the CLAMP device grippers and these wider objects (wider objects led to inconsistent and partial surface area contact between sensors and object surfaces). This was not the case during human data collection because humans used the LED signals to close the loop on their grasping policy such that the contact was complete and consistent. To test our hypothesis, we finetune the CLAMP model on 30% data out of all data composed of only objects grasped by all embodiments. With 0.96 test accuracy, “Franka with CLAMP” slightly outperforms “Franka with Parallel Jaw” (0.95) and “WidowX with Parallel Jaw” (0.90), which better reflects the effect of device and various robot embodiment differences.
D.3 Sorting recyclable from non-recyclable items
@@
D.4 Retrieving objects from a cluttered bag
For this experiment, we place three objects in a cloth bag: a T-shirt, a steel fork, and a bag of popcorn. The robot captures an image, reaches into the bag, grasps each object, and retrieves objects if the CLAMP model predicts it to be one of (aluminium, brass, or steel). We use a total of two object sequences across the 13 trials: fork $\rightarrow$ T-shirt $\rightarrow$ popcorn and T-shirt $\rightarrow$ fork $\rightarrow$ popcorn. We observe that grasping the bag first or second causes the cloth bag to crumple and thus prevent further grasps.
D.5 Separating ripe from overripe bananas